OpenStudy (anonymous):

For the sets A = (1, 2], B = [−2, 1] and U=Reals, there's some identities I have to find for them I'm not familiar with. A^c , B^c , A ∩ B, A ∪ B, A\B, B \A, and A delta B

OpenStudy (anonymous):

\[A^c=\{x|x\notin A\}\]so in this case it would be \[(-\infty, 1]\cup (2,\infty)\]

OpenStudy (anonymous):

is there a list of these identity functions somewhere? I think I could figure it out if I had something to refer to, but I don't unfortunately

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

like, I'm not sure what A^c or A delta B or A \ B even mean

OpenStudy (anonymous):

A^c means everything not in A

OpenStudy (anonymous):

A\B means everything that is in A but not in B

OpenStudy (anonymous):

and i am not sure about the delta notation, but my guess it it means everything in A or B but not both

OpenStudy (anonymous):

Well, I guessed it meant delta because the symbol used is a small triangle.

OpenStudy (anonymous):

aks \[A\Delta B= (A\cup B)-(A\cap B)\]

OpenStudy (anonymous):

not standard notation, usually you see \[A\oplus B\]

OpenStudy (anonymous):

you ok with unions and intersections?

OpenStudy (anonymous):

so the big U is the union of set A and B, meaning all their values together? and the big n is the intersection, only what they have in common?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

http://www.youtube.com/watch?v=En8fI2ixepo

OpenStudy (anonymous):

thanks a lot man!

OpenStudy (anonymous):

yw